Find f'(x), the derivative of f at X, for F(X) = 1/x
Answer by stanbon(75887)

Find f'(x), the derivative of f at X, for F(X) = 1/x
F'(X) = lim/h>0 f(x+H)- f(x)/h
[f(x+h) - f(x)]/h = [1/(x+h) - 1/x] / h
------
=[x - (x+h)]/[x(x+h)]/h = [-h]/[x*h(x+h)] = [-1/[x(x+h)]
----
Take the limit as h goes to 0 to get: -1/x^2
============================
Ans:: f'(F(x)) = -1/x^2
-----------------------
Cheers,
Stan H.
-------------